# GenMol: A Drug Discovery Generalist with Discrete Diffusion

### v1

@@ 5 Experiments
5.1 De Novo Generation
Setup.
In de novo generation, the goal is to generate valid, unique, and diverse molecules. We generate 1,000 molecules and evaluate them with the following metrics, following Noutahi et al. (2024). Validity is the fraction of generated molecules that are chemically valid. Uniqueness is the fraction of valid molecules that are unique. Diversity is defined as the average pairwise Tanimoto distance between the Morgan fingerprints of the generated molecules. We further introduce quality, the fraction of generated molecules that are valid, unique, drug-like, and synthesizable, to provide a single metric that summarizes validity and uniqueness and evaluates the model’s ability to generate chemically reasonable molecules. Here, drug-like and synthesizable molecules are defined as those satisfying quantitative estimate of drug-likeness (QED) (Bickerton et al., 2012) $\geq 0.6$ and synthetic accessibility (SA) (Ertl & Schuffenhauer, 2009) $\leq 4$, respectively, following Jin et al. (2020). Further details are provided in Section A.3.
Results.
The results of de novo generation are shown in Table 4. Both GenMol and SAFE-GPT (Noutahi et al., 2024) show near-perfect uniqueness, while GenMol significantly outperforms SAFE-GPT in terms of validity, quality, and sampling time. Thanks to the non-autoregressive parallel decoding scheme, GenMol can predict multiple tokens simultaneously and shows much faster sampling as $N$, the number of tokens to unmask at each time step, increases. Notably, GenMol with $N=3$ shows higher quality than SAFE-GPT with 2.7x shorter sampling time and similar diversity. Furthermore, GenMol has the ability to balance between quality and diversity by adjusting the values of the softmax temperature $\tau$ and the randomness $r$ of the confidence sampling described in Section 4.1. This balance is also shown in Figure 4, demonstrating that GenMol generates molecules along the Pareto frontier of the quality-diversity trade-off. Further analysis on quality and diversity are provided in Section B.1 and Section B.2.
5.2 Fragment-constrained Generation
Setup.
In fragment-constrained generation, the goal is to complete molecules given a set of fragments, a frequently encountered setting in real-world drug discovery scenarios. We use the benchmark proposed by Noutahi et al. (2024), which uses input fragments extracted from 10 known drugs to perform the linker design, scaffold morphing, motif extension, scaffold decoration, and superstructure generation tasks. In addition to validity, uniqueness, diversity, and quality introduced in Section 5.1, distance, the average Tanimoto distance between the generated molecules and the original molecules, is also measured. 100 molecules are generated for each drug and averaged to calculate the metrics. Further details are provided in Section A.4.
Results.
The results of fragment-constrained generation are shown in Table 2. GenMol significantly outperforms SAFE-GPT on most metrics across the tasks, demonstrating its general applicability to a variety of fragment constraint generation settings. Especially, GenMol is superior to SAFE-GPT in generating high-quality molecules while preserving high diversity under the fragment constraints, again validating the superiority of GenMol in balancing between quality and diversity.
5.3 Goal-directed Hit Generation
Table 3: Goal-directed hit generation results. The results are the means of PMO AUC top-10 of 3 runs. The results for $f$-RAG (Lee et al., 2024a), Genetic GFN (Kim et al., 2024) and Mol GA (Tripp & Hernández-Lobato, 2023) are taken from the respective papers and the results for REINVENT and Graph GA are taken from Gao et al. (2022). The best results are highlighted in bold.
  Oracle GenMol GenMol-TR GenMol-NR $f$-RAG Genetic GFN Mol GA REINVENT Graph GA
albuterol_similarity 0.932 0.895 0.872 0.977 0.949 0.896 0.882 0.838
amlodipine_mpo 0.804 0.802 0.769 0.749 0.761 0.688 0.635 0.661
celecoxib_rediscovery 0.826 0.821 0.859 0.778 0.802 0.567 0.713 0.630
deco_hop 0.953 0.945 0.917 0.936 0.733 0.649 0.666 0.619
drd2 0.995 0.995 0.995 0.992 0.974 0.936 0.945 0.964
fexofenadine_mpo 0.894 0.886 0.875 0.856 0.856 0.825 0.784 0.760
gsk3b 0.986 0.985 0.985 0.969 0.881 0.843 0.865 0.788
isomers_c7h8n2o2 0.934 0.934 0.897 0.955 0.969 0.878 0.852 0.862
isomers_c9h10n2o2pf2cl 0.833 0.830 0.816 0.850 0.897 0.865 0.642 0.719
jnk3 0.856 0.848 0.845 0.904 0.764 0.702 0.783 0.553
median1 0.397 0.397 0.397 0.340 0.379 0.257 0.356 0.294
median2 0.355 0.350 0.349 0.323 0.294 0.301 0.276 0.273
mestranol_similarity 0.981 0.980 0.970 0.671 0.708 0.591 0.618 0.579
osimertinib_mpo 0.876 0.876 0.876 0.866 0.860 0.844 0.837 0.831
perindopril_mpo 0.703 0.703 0.697 0.681 0.595 0.547 0.537 0.538
qed 0.942 0.942 0.927 0.939 0.942 0.941 0.941 0.940
ranolazine_mpo 0.821 0.818 0.809 0.820 0.819 0.804 0.760 0.728
scaffold_hop 0.628 0.621 0.617 0.576 0.615 0.527 0.560 0.517
sitagliptin_mpo 0.573 0.560 0.573 0.601 0.634 0.582 0.021 0.433
thiothixene_rediscovery 0.687 0.686 0.650 0.584 0.583 0.519 0.534 0.479
troglitazone_rediscovery 0.867 0.853 0.801 0.448 0.511 0.427 0.441 0.390
valsartan_smarts 0.797 0.797 0.739 0.627 0.135 0.000 0.178 0.000
zaleplon_mpo 0.569 0.569 0.406 0.486 0.552 0.519 0.358 0.346
Sum 18.208 18.091 17.641 16.928 16.213 14.708 14.196 13.751
Setup.
In goal-directed hit generation, the goal is to generate hits, i.e., molecules that are optimized in terms of target chemical properties. Following Lee et al. (2024b) and Lee et al. (2024a), we first construct a fragment vocabulary by fragmenting the molecules in the ZINC250k dataset (Irwin et al., 2012). During generation, a molecule is generated by randomly selecting two fragments from the vocabulary and concatenating them, and then applying the fragment remasking using GenMol. The fragment vocabulary is dynamically updated throughout the generation process, allowing exploration beyond the initial fragments. We adopt the practical molecular optimization (PMO) benchmark (Gao et al., 2022) which contains 23 optimization tasks. Following the benchmark setting, the maximum number of oracle calls is set to 10,000 and optimization performance is measured using the area under the curve (AUC) of the average top-10 property scores versus oracle calls. Further details are provided in Section A.5.
Baselines.
As our baselines, we adopt the three recent state-of-the-art methods of the PMO benchmark. We also report performance of the top two methods reported by Gao et al. (2022). Note that as the paper of Gao et al. (2022) contains the results of a total of 25 methods, comparing GenMol to the top methods is equivalent to comparing it to 25 methods. The results of additional baselines including the standard deviations are provided in Table 5, Table 6, and Table 7. Graph GA (Jensen, 2019) is a GA that performs a random fragment-level crossover and a random atom- or bond-level mutation, and $f$-RAG (Lee et al., 2024a) is a method that uses fragment-level retrieval-augmented generation (RAG) and Graph GA-based fragment modification. Genetic GFN (Kim et al., 2024) utilizes Graph GA to guide a GFlowNet. Mol GA (Tripp & Hernández-Lobato, 2023) is a tuned version of Graph GA. REINVENT (Olivecrona et al., 2017) is a reinforcement learning (RL) model based on the SMILES representation. We additionally compare GenMol with GenMol-TR and GenMol-NR, the ablated variants of GenMol which will be discussed in Section 5.5.
Results.
The results of goal-directed hit generation are shown in Table 3. GenMol significantly outperforms the previous methods in terms of the sum AUC top-10 value and exhibits the best performance in 18 out of 23 tasks by a large margin. These results verify that the proposed optimization strategy of fragment remasking is effective in exploring chemical space and discovering optimized hit molecules. The results of GenMol trained on the ZINC250k dataset are shown in Table 8.
5.4 Goal-directed Lead Optimization
Table 4: Lead optimization results (kcal/mol). The results are the docking scores of the most optimized leads. Lower is better, and the best results are highlighted in bold.
  Target protein Seed score $\delta=0.4$ $\delta=0.6$
12
GenMol GenMol-TR GenMol-NR RetMol Graph GA GenMol GenMol-TR GenMol-NR RetMol Graph GA
-7.3 -10.6 -8.7 -8.2 -9.0 -8.3 -10.4 -8.1 -8.3 - -8.6
parp1 -7.8 -11.0 -8.2 -8.3 -10.7 -8.9 -10.5 -8.2 -8.1 - -8.1
-8.2 -11.5 -10.9 - - - -9.5 -8.7 - - -
-6.4 -7.5 -7.8 -7.2 -8.0 -7.8 -7.0 - - -7.6 -7.6
fa7 -6.7 -8.4 -8.0 -8.3 - -8.2 -7.6 -7.6 - - -7.6
-8.5 - - - - - - - - - -
-4.5 -12.9 -12.8 - -12.1 -11.7 -12.1 -11.0 - - -11.3
5ht1b -7.6 -12.3 -11.8 -11.9 -9.0 -12.1 -12.2 -11.8 -12.6 -10.0 -12.0
-9.8 -11.8 -11.3 - - - -11.0 -10.7 - - -
-9.3 -10.8 -10.7 -9.7 - -9.8 - - - - -
braf -9.4 -10.2 -10.8 - -11.6 - -10.2 - -9.6 - -
-9.8 -10.6 -10.5 - - -11.6 -10.2 -10.0 - - -10.4
-7.7 -10.0 -8.7 -8.6 -8.2 -8.7 -9.5 -8.1 -8.0 - -8.1
jak2 -8.0 -9.9 -9.0 -8.9 -9.0 -9.2 -9.4 -9.0 -8.9 - -9.2
-8.6 -9.5 -9.2 - - - - - - - -
Setup.
In goal-directed lead optimization, the goal is to generate leads when an initial seed molecule is given. Here, leads are defined as the molecules that exhibit improved target properties while maintaining the similarity with the given seed. We adopt the setting of Wang et al. (2023), where the objective is to optimize the binding affinity to the target protein while satisfying the following constraints: $\text{QED}\geq 0.6$, $\text{SA}\leq 4$, and $\text{sim}\geq\delta$ where $\delta\in\{0.4,0.6\}$ and sim is the pairwise Tanimoto similarity between the Morgan fingerprints of the generated molecules and the seed. Optimization performance is evaluated by the docking score of the most optimized molecule that satisfies the above constraints. Following Lee et al. (2023), we adopt five target proteins, parp1, fa7, 5ht1b, braf, and jak2 to avoid bias in target selection. For each target, three molecules from its known active molecules are selected and each is given as a seed molecule, yielding a total of 15 tasks. As in Section 5.2, a fragment vocabulary is first constructed by fragmenting the seed molecule. A molecule is generated by randomly attaching two fragments from the vocabulary, and then applying GenMol’s fragment remasking. If the generated molecule is lead, its fragments are added to the fragment vocabulary. Further details are provided in Section A.6.
Baselines.
Following Wang et al. (2023), we adopt Graph GA introduced in Section 5.3 and RetMol (Wang et al., 2023), a retrieval-based molecular optimization method that uses retrieved molecules to guide the generation of a molecular language model, as our baselines. GenMol-TR and GenMol-NR are the ablated variants of GenMol which will be discussed in Section 5.5.
Results.
The results of goal-directed lead optimization are shown in Table 4. As shown in the table, GenMol outperforms the baselines in most tasks. Note that baselines often fail, i.e., they cannot generate molecules with a higher binding affinity than the seed molecule while satisfying the constraints, especially under the harsher ($\delta=0.6$) similarity constraint. In contrast, GenMol is able to successfully optimize most seed molecules, validating its effectiveness in exploring chemical space to optimize given molecules and discover promising lead molecules. The results of GenMol trained on the ZINC250k dataset (Irwin et al., 2012) are shown in Table 9.
5.5 Ablation Study
To examine the effect of the proposed fragment remasking scheme of GenMol, we compare GenMol with GenMol-token remasking (GenMol-TR), GenMol that randomly masks individual tokens instead of masking a fragment chunk, and GenMol-no remasking (GenMol-NR), GenMol that performs no remasking at all (i.e., relies solely on attaching random fragments of the fragment vocabulary to generate molecules), in Table 3 and Table 4. Both GenMol and GenMol-token remasking outperform GenMol-no remasking. Especially, GenMol-no remasking fails more frequently in the lead optimization task, highlighting the importance of the exploration strategy through remasking of masked diffusion models. On the other hand, GenMol outperforms GenMol-token remasking, proving that the proposed fragment remasking strategy is well aligned with chemical intuition and is therefore effective in exploring the chemical space to discover novel chemical optima.
6 Conclusion
We proposed GenMol, a versatile molecule generation framework designed to deal with various scenarios of drug discovery. By integrating discrete diffusion with the SAFE molecular representation, GenMol is able to effectively and efficiently generate molecules under settings that simulate a variety of drug discovery problems. Especially, the fragment remasking strategy that uses fragments as an explorative unit allows GenMol to effectively explore chemical space and discover chemical optima. Our experimental results showed that GenMol is capable of achieving state-of-the-art results in a wide range of drug discovery tasks, demonstrating the potential of GenMol as a unified and versatile tool for drug discovery.
Appendix A Experimental Details
A.1 Computing Resources
GenMol was trained using 8 NVIDIA A100 GPUs. The training took approximately 6 hours. All the molecular generation experiments were conducted using a single NVIDIA A100 GPU and 32 CPU cores.
A.2 Training GenMol
Figure 5: Illustration of the fragment-constrained motif extension process of (a) GenMol and (b) autoregressive models. GenMol starts by sampling the length of the sequence and then filling the sequence with mask tokens corresponding to the sampled length. GenMol employs parallel decoding where all tokens are decoded simultaneously under discrete diffusion, and confirms only the most confident predictions. The decoding proceeds progressively until all mask tokens are predicted. In contrast, autoregressive models like SAFE-GPT (Noutahi et al., 2024) need to predict one token per step, requiring many more decoding steps.
In this section, we describe the details for training GenMol. We used the BERT (Devlin et al., 2019) architecture of the HuggingFace Transformers library (Wolf et al., 2019) with the default configuration, except that we set max_position_embeddings to 256. We used the SAFE dataset and SAFE tokenizer (Noutahi et al., 2024) that has a vocabulary size of $K=1880$. We set the batch size to 2048, the learning rate to $43e-$ 4, and the number of training steps to 50k. We used the log-linear noise schedule of Sahoo et al. (2024) and the AdamW optimizer (Loshchilov & Hutter, 2019) with $\beta_{1}=0.9$ and $\beta_{2}=0.999$.
A.3 De Novo Generation
A.4 Fragment-constrained Generation
We used $N=1$. We performed the grid search with the search space $\tau\in\{0.5,0.8,1,1.2,1.5\}$ and $r\in\{1,1.2,2,3\}$, and set $r$ to 3 for linker design and scaffold morphing, 1.2 for motif extension, and 2 for scaffold decoration and superstructure generation. We set $\tau$ to 1.2 for all the tasks. The lengths of the mask chunks were sampled from the ZINC250k distribution.
A.5 Goal-directed Hit Generation
In this section, we describe the details for conducting experiments in Section 5.3. To construct an initial fragment vocabulary, we adopted a simple decomposition rule that randomly cut one of the non-ring single bonds of a given molecules three times and apply it to the ZINC250k dataset. We set the size of the fragment vocabulary to 100. We applied the warmup scheme that let GenMol generate molecules by concatenating two randomly chosen fragments without fragment remasking for the first 1,000 generations. We used $N=1$, $\tau=1.2$, and $r=2$.
A.6 Goal-directed Lead Optimization
In this section, we describe the details for conducting experiments in Section 5.4. Following the setting of Wang et al. (2023), for each target protein and each seed molecule, we run 10 optimization iterations with 100 generation per iteration. We adopted the same decomposition rule described in Section A.5 and used $N=1$, $\tau=1.2$, and $r=2$.
Appendix B Additional Experimental Results
B.1 QED and SA Distributions in De Novo Generation
B.2 Analysis on Quality and Diversity in De Novo Generation
Figure 6: QED and SA distributions of molecules in de novo generation.
B.3 Full Goal-directed Hit Generation Results
We provide the full results of Table 3 including the additional baselines from Gao et al. (2022) and the standard deviations in Table 5, Table 6, and Table 7. As shown in the tables, GenMol outperforms all the 28 baselines by a large margin.
@@ B.3 Full Goal-directed Hit Generation Results
  Oracle GenMol $f$-RAG Genetic GFN Mol GA
albuterol_similarity 0.932 $\pm$ 0.007 0.977 $\pm$ 0.002 0.949 $\pm$ 0.010 0.896 $\pm$ 0.035
amlodipine_mpo 0.804 $\pm$ 0.006 0.749 $\pm$ 0.019 0.761 $\pm$ 0.019 0.688 $\pm$ 0.039
celecoxib_rediscovery 0.826 $\pm$ 0.018 0.778 $\pm$ 0.007 0.802 $\pm$ 0.029 0.567 $\pm$ 0.083
deco_hop 0.953 $\pm$ 0.016 0.936 $\pm$ 0.011 0.733 $\pm$ 0.109 0.649 $\pm$ 0.025
drd2 0.995 $\pm$ 0.000 0.992 $\pm$ 0.000 0.974 $\pm$ 0.006 0.936 $\pm$ 0.016
fexofenadine_mpo 0.894 $\pm$ 0.028 0.856 $\pm$ 0.016 0.856 $\pm$ 0.039 0.825 $\pm$ 0.019
gsk3b 0.986 $\pm$ 0.003 0.969 $\pm$ 0.003 0.881 $\pm$ 0.042 0.843 $\pm$ 0.039
isomers_c7h8n2o2 0.934 $\pm$ 0.002 0.955 $\pm$ 0.008 0.969 $\pm$ 0.003 0.878 $\pm$ 0.026
isomers_c9h10n2o2pf2cl 0.833 $\pm$ 0.014 0.850 $\pm$ 0.005 0.897 $\pm$ 0.007 0.865 $\pm$ 0.012
jnk3 0.856 $\pm$ 0.016 0.904 $\pm$ 0.004 0.764 $\pm$ 0.069 0.702 $\pm$ 0.123
median1 0.397 $\pm$ 0.000 0.340 $\pm$ 0.007 0.379 $\pm$ 0.010 0.257 $\pm$ 0.009
median2 0.355 $\pm$ 0.003 0.323 $\pm$ 0.005 0.294 $\pm$ 0.007 0.301 $\pm$ 0.021
mestranol_similarity 0.981 $\pm$ 0.003 0.671 $\pm$ 0.021 0.708 $\pm$ 0.057 0.591 $\pm$ 0.053
osimertinib_mpo 0.876 $\pm$ 0.008 0.866 $\pm$ 0.009 0.860 $\pm$ 0.008 0.844 $\pm$ 0.015
perindopril_mpo 0.703 $\pm$ 0.009 0.681 $\pm$ 0.017 0.595 $\pm$ 0.014 0.547 $\pm$ 0.022
qed 0.942 $\pm$ 0.000 0.939 $\pm$ 0.001 0.942 $\pm$ 0.000 0.941 $\pm$ 0.001
ranolazine_mpo 0.821 $\pm$ 0.011 0.820 $\pm$ 0.016 0.819 $\pm$ 0.018 0.804 $\pm$ 0.011
scaffold_hop 0.628 $\pm$ 0.008 0.576 $\pm$ 0.014 0.615 $\pm$ 0.100 0.527 $\pm$ 0.025
sitagliptin_mpo 0.573 $\pm$ 0.050 0.601 $\pm$ 0.011 0.634 $\pm$ 0.039 0.582 $\pm$ 0.040
thiothixene_rediscovery 0.687 $\pm$ 0.125 0.584 $\pm$ 0.009 0.583 $\pm$ 0.034 0.519 $\pm$ 0.041
troglitazone_rediscovery 0.867 $\pm$ 0.022 0.448 $\pm$ 0.017 0.511 $\pm$ 0.054 0.427 $\pm$ 0.031
valsartan_smarts 0.797 $\pm$ 0.033 0.627 $\pm$ 0.058 0.135 $\pm$ 0.271 $0$$\pm$ 0.000
zaleplon_mpo 0.569 $\pm$ 0.005 0.486 $\pm$ 0.004 0.552 $\pm$ 0.033 0.519 $\pm$ 0.029
Sum 18.208 16.928 16.213 14.708
@@
B.4 Goal-directed Hit Generation Results of GenMol Variants
We provide the full results of the internal baselines in Table 3, GenMol-TR and GenMol-NR, and GenMol trained on the ZINC250k dataset (Irwin et al., 2012) instead of the SAFE dataset, GenMol-ZINC250k, in Table 8. As shown in the table, GenMol-ZINC250k shows comparable results to GenMol, while GenMol-TR and GenMol-NR show inferior performance to GenMol as discussed in Section 5.5.

### v2

@@ E.6 Ablation Study on MCG in Fragment-constrained Generation
GenMol Linker design 100.0 $\pm$ 0.0 (+0.0) 83.7 $\pm$ 0.5 (+0.0) 21.9 $\pm$ 0.4 (+0.0) 0.547 $\pm$ 0.002 (+0.000) 0.563 $\pm$ 0.003 (+0.000)
Scaffold morphing 100.0 $\pm$ 0.0 (+0.0) 83.7 $\pm$ 0.5 (+0.0) 21.9 $\pm$ 0.4 (+0.0) 0.547 $\pm$ 0.002 (+0.000) 0.563 $\pm$ 0.003 (+0.000)
Motif extension 082.9 $\pm$ 0.1 (+5.7) 77.5 $\pm$ 0.1 (-0.3) 30.1 $\pm$ 0.4 (+2.6) 0.617 $\pm$ 0.002 (+0.000) 0.682 $\pm$ 0.001 (+0.000)
Scaffold decoration 096.6 $\pm$ 0.8 (-0.2) 82.7 $\pm$ 1.8 (+4.7) 31.8 $\pm$ 0.5 (+2.2) 0.591 $\pm$ 0.001 (+0.015) 0.651 $\pm$ 0.001 (+0.001)
Superstructure generation 097.5 $\pm$ 0.9 (-0.7) 83.6 $\pm$ 1.0 (+5.3) 34.8 $\pm$ 1.0 (+1.5) 0.599 $\pm$ 0.009 (+0.025) 0.762 $\pm$ 0.007 (+0.005)
E.7 Full Goal-directed Hit Generation Results
We provide the full results of Table 3 including the additional baselines from Gao et al. (2022) in Table 13, Table 14, and Table 15. As shown in the tables, GenMol outperforms all baselines by a large margin.
On the other hand, to simulate the fragment-based drug discovery (FBDD) scenario following Lee et al. (2024a), we have assumed that a high-quality fragment vocabulary of size 100 is given for each task in Table 3. Here, similar to Wang et al. (2023), we also provide the results of GenMol (10 fragments), which starts with only 10 of these fragments to simulate a scenario where high-quality fragments are sparse, in Table 13.
E.8 Full Goal-directed Hit Generation Results in Ablation Study
We provide the full results of the ablated GenMol variants baselines in the goal-directed hit generatino task in Table 5 in Table 16. As shown in the table, A, A + Token remasking, A + GPT remasking and A + Fragment remasking show inferior performance to A + F + MCG, the full GenMol, as discussed in Section 5.5.
E.9 Ablation Study in Goal-directed Lead Optimization
We compare the results of the ablated GenMol baselines in the goal-directed lead optimization task in Table 17. As in Table 5, A + Token remasking, A + GPT remasking and A + Fragment remasking all outperform A, demonstrating the importance of further modification on top of A through remasking. Moreover, A + Fragment remasking generally outperforms A + Token remasking and A + GPT remasking, showing the effectiveness of the proposed remasking strategy that uses fragments as the exploration unit under discrete diffusion. Lastly, A + F + MCG shows further improved performance compared to A + F, validating the effectiveness of MCG in the lead optimization task.
E.10 Examples of Generated Molecules
We provide examples of the molecules generated by GenMol ($N=1$, $\tau=0.5$, $r=0.5$) on de novo generation Figure 7. We provide examples of generated molecules on fragment-constrained generation in Figure 8.
@@ E.10 Examples of Generated Molecules
  Oracle GenMol GenMol (10 fragments) $f$-RAG Genetic GFN Mol GA
albuterol_similarity 0.937 $\pm$ 0.010 0.847 $\pm$ 0.036 0.977 $\pm$ 0.002 0.949 $\pm$ 0.010 0.896 $\pm$ 0.035
amlodipine_mpo 0.810 $\pm$ 0.012 0.762 $\pm$ 0.012 0.749 $\pm$ 0.019 0.761 $\pm$ 0.019 0.688 $\pm$ 0.039
celecoxib_rediscovery 0.826 $\pm$ 0.018 0.619 $\pm$ 0.005 0.778 $\pm$ 0.007 0.802 $\pm$ 0.029 0.567 $\pm$ 0.083
deco_hop 0.960 $\pm$ 0.010 0.957 $\pm$ 0.016 0.936 $\pm$ 0.011 0.733 $\pm$ 0.109 0.649 $\pm$ 0.025
drd2 0.995 $\pm$ 0.000 0.995 $\pm$ 0.000 0.992 $\pm$ 0.000 0.974 $\pm$ 0.006 0.936 $\pm$ 0.016
fexofenadine_mpo 0.894 $\pm$ 0.028 0.806 $\pm$ 0.008 0.856 $\pm$ 0.016 0.856 $\pm$ 0.039 0.825 $\pm$ 0.019
gsk3b 0.986 $\pm$ 0.003 0.985 $\pm$ 0.004 0.969 $\pm$ 0.003 0.881 $\pm$ 0.042 0.843 $\pm$ 0.039
isomers_c7h8n2o2 0.942 $\pm$ 0.004 0.984 $\pm$ 0.002 0.955 $\pm$ 0.008 0.969 $\pm$ 0.003 0.878 $\pm$ 0.026
isomers_c9h10n2o2pf2cl 0.833 $\pm$ 0.014 0.866 $\pm$ 0.010 0.850 $\pm$ 0.005 0.897 $\pm$ 0.007 0.865 $\pm$ 0.012
jnk3 0.906 $\pm$ 0.023 0.828 $\pm$ 0.007 0.904 $\pm$ 0.004 0.764 $\pm$ 0.069 0.702 $\pm$ 0.123
median1 0.398 $\pm$ 0.000 0.336 $\pm$ 0.008 0.340 $\pm$ 0.007 0.379 $\pm$ 0.010 0.257 $\pm$ 0.009
median2 0.359 $\pm$ 0.004 0.354 $\pm$ 0.000 0.323 $\pm$ 0.005 0.294 $\pm$ 0.007 0.301 $\pm$ 0.021
mestranol_similarity 0.982 $\pm$ 0.000 0.991 $\pm$ 0.002 0.671 $\pm$ 0.021 0.708 $\pm$ 0.057 0.591 $\pm$ 0.053
osimertinib_mpo 0.876 $\pm$ 0.008 0.870 $\pm$ 0.004 0.866 $\pm$ 0.009 0.860 $\pm$ 0.008 0.844 $\pm$ 0.015
perindopril_mpo 0.718 $\pm$ 0.012 0.695 $\pm$ 0.004 0.681 $\pm$ 0.017 0.595 $\pm$ 0.014 0.547 $\pm$ 0.022
qed 0.942 $\pm$ 0.000 0.943 $\pm$ 0.000 0.939 $\pm$ 0.001 0.942 $\pm$ 0.000 0.941 $\pm$ 0.001
ranolazine_mpo 0.821 $\pm$ 0.011 0.777 $\pm$ 0.016 0.820 $\pm$ 0.016 0.819 $\pm$ 0.018 0.804 $\pm$ 0.011
scaffold_hop 0.628 $\pm$ 0.008 0.648 $\pm$ 0.005 0.576 $\pm$ 0.014 0.615 $\pm$ 0.100 0.527 $\pm$ 0.025
sitagliptin_mpo 0.584 $\pm$ 0.034 0.588 $\pm$ 0.064 0.601 $\pm$ 0.011 0.634 $\pm$ 0.039 0.582 $\pm$ 0.040
thiothixene_rediscovery 0.692 $\pm$ 0.123 0.569 $\pm$ 0.013 0.584 $\pm$ 0.009 0.583 $\pm$ 0.034 0.519 $\pm$ 0.041
troglitazone_rediscovery 0.867 $\pm$ 0.022 0.848 $\pm$ 0.040 0.448 $\pm$ 0.017 0.511 $\pm$ 0.054 0.427 $\pm$ 0.031
valsartan_smarts 0.822 $\pm$ 0.042 0.803 $\pm$ 0.011 0.627 $\pm$ 0.058 0.135 $\pm$ 0.271 $0$$\pm$ 0.000
zaleplon_mpo 0.584 $\pm$ 0.011 0.571 $\pm$ 0.016 0.486 $\pm$ 0.004 0.552 $\pm$ 0.033 0.519 $\pm$ 0.029
Sum 18.362 17.643 16.928 16.213 14.708